In a single slit diffraction experiment, the screen is placed 2.45m away from the slit. Light with a wavelength of 685nm forms a pattern on the screen and the fourteenth dark fringe form 76.0cm above center of the pattern. What is the width of the slit?
Answer #1

for dark fringe dyn = (2n-1) D for n=14 dyn = [2(14)-1] 272 dyny D d 277D 2414 27*685* 09* 2.45 2*0.76 d=29.81 * 107 -6 -slit